4 u bende bekiyorum Anne saliyorum onun dişinda he foll haa yukluyorum felan Derin Ipekgil tion MacBook Air Question 3 (15 pts): As a famous music producer, you ha has a very good voice and her singing is blessed. Moreover, she is an attractive lady and a great dancer, which are great attributes for a singer. She has sung in night clubs and for special events like new years and weddings. She worked only for 3 months last year and her annual income from singing activities reached to $30,000. You believe that if she works full year ar would be equal to the amount in the following table for the next 5 years. ound, her income Year Future Income S 200,000 S 400,000 S 450,000 S 300,000 $ 50,000 Besides, she is a composer and she has already sold some compositions of hers to other singers. She made S150.000 from her compositions only last year. Because she is very productive, you estimate she will increase her income 5% per year for the next 5 years. The problem with her is that, she needs to stay in her mountain house for a very long time to compose new songs. Her seclusion period can vary between 6 months and 1 year, during which she secludes herself in and refuses to make contact to others. Therefore, she cannot perform on the stage during these periods. Another issue about the agreement is that she requires $1,000,000 if she is recruited as a singer and S500,000 if she is recruited as a composer by the time of the agrcement. So, you can use her only as a singer or only as a composer. To make a final decision, please answer following questions, assuming discount rate 10%. a) (2 pts) Please fill in the gaps in the table (a.b,c,d,e,f.g) Year Singer Composer (500,000) 200,000 400,000 300,000 50,000

b) if you apply "payback criterion" which option would you choose? Singer or composer? (if you use excel to solve this question, please copy your solution)

c) if you apply "NPV" criterion, which option would you choose? why? (if you use excel to solve this question, please copy your solution and explain it verbally)

d) If you apply "IRR" criterion, which investment you choose? why? (if you use excel to solve this question, please copy your solution and explain it verbally)

e) is there a conflict or agreement between the payback, NPV and IRR rules) How do you interpret this situation?

f) what is your final decision? Why?
Answer #1

Part a).

Year Singer Composer
0 -1,000,000.00 -500,000.00
1 200,000.00 157,500.00
2 400,000.00 165,375.00
3 450,000.00 173,643.75
4 300,000.00 182,325.94
5 50,000.00 191,442.23

a = -1,000,000 (since that the the signing amount for Cindy as a singer)

Her income as a composer is expected to increase by 5% per year for the next 5 years when her current earnings as a composer are 150,000 so

b = Year 1 earnings = 1.05*150,000 = 157,500

c = Year 2 earnings = 1.05*157,500 = 163,375

d = Year 3 earnings = 1.05*163,375 = 173,643.75

e = Year 4 earnings = 1.05*173,643.75 = 182,325.94

f = Year 5 earnings = 1.05*182,325.94 = 191,442.23

Part b).

Payback period (as a Singer)

Year 0 1 2 3 4 5
Singer -1,000,000.00 200,000.00 400,000.00 450,000.00 300,000.00 50,000.00
Cumulative cash flow -1,000,000.00 -800,000.00 -400,000.00 50,000.00 350,000.00 400,000.00

We see that the cumulative cash flow turns from negative to positive in Year 3. The amount of cash flow which is needed to be turn the cash flow positive is the cumulative cash flow in Year 2 which is -400,000

The fraction of Year 3 = Cumulative cash flow in Year 2/Cash flow in Year 3 = 400,000/450,000 = 0.89

Thus, payback period = 2 years + 0.89 year = 2.89 years. (Answer)

Payback period (as a Composer)

Year 0 1 2 3 4 5
Composer -500,000.00 157,500.00 165,375.00 173,643.75 182,325.94 191,442.23
Cumulative cash flow -500,000.00 -342,500.00 -177,125.00 -3,481.25 178,844.69 370,286.92

We see that the cumulative cash flow turns from negative to positive in Year 4. The amount of cash flow which is needed to be turn the cash flow positive is the cumulative cash flow in Year 3 which is -3,481.25

The fraction of Year 4 = Cumulative cash flow in Year 3/Cash flow in Year 4 = 3,481.25/182,325.94 = 0.019

Thus, payback period = 3 years + 0.019 year = 3.019 years. (Answer)

Part c). NPV (as a Singer) = 86,438.46

NPV (as a Composer) = 153,718.13

(Calculations shown below)

Part d). IRR (as a Singer) = 13.58%

IRR (as a Composer) = 21.00%

(IRR calculated using the cash flows for each option and the IRR() function in excel)

Formula Year (n) 0 1 2 3 4 5
Singer cash flow -1,000,000.00 200,000.00 400,000.00 450,000.00 300,000.00 50,000.00
1/(1+d)^n Discount Factor @10% 1.000 0.909 0.826 0.751 0.683 0.621
Cash flow*Discount factor Present Value of cash flow -1,000,000.00 181,818.18 330,578.51 338,091.66 204,904.04 31,046.07
Sum of all PVs of cash flow NPV (Singer) 86,438.46
IRR (Singer) 13.58%
Formula Year (n) 0 1 2 3 4 5
Composer cash flow -500,000.00 157,500.00 165,375.00 173,643.75 182,325.94 191,442.23
1/(1+d)^n Discount Factor @10% 1.000 0.909 0.826 0.751 0.683 0.621
Cash flow*Discount factor Present Value of cash flow -500,000.00 143,181.82 136,673.55 130,461.12 124,531.07 118,870.57
Sum of all PVs of cash flow NPV (Composer) 153,718.13
IRR (Composer) 21.00%

Part e). There is no conflict between the NPV and IRR rules since both indicate that the agreement as a Composer should be chosen as it gives a higher positive NPV and a greater IRR. However, if we consider the Payback Period then the option as a a Singer should be chosen since it has a shorter payback period. This happens because the Payback Period method does not take the Present Value of cash flows into account and also because the cash inflows for the Singer option are much better than that for the Composer option (except in the last year).

Part f). NPV is always a better criterion to use for deciding whether to undertake a project so the option as a Composer should be chosen as it has a much better NPV compared to the NPV as a Singer.